In a race, Car A starts 1 mile behind Car B. Car A is traveling at 65 miles per hour, while Car B is traveling at 45 miles per hour. How long will it take for Car A to overtake Car B?
Answer by reynard2007(52)


The moment Car A overtakes Car B is when they have travelled the same distance. However, Car A has to make up for an extra mile since he started 1 mile behind. Let x be the number of hours.
 
Distance = Rate * Time
Distance of Car A = Distance of Car B + 1
 
(65 * x) = (45 * x ) + 1
65x-45x = 1
20x = 1
x = 1/20 hour = 3 minutes
 
Car A will overtake Car B in 3 minutes